OpenStudy (anonymous):

Solve for y. y/4 - (y-1)/2 (> or equal to) 3

OpenStudy (anonymous):

\[\frac{y}{4}-\frac{y-1}{2}\geq 3\] multiply by 4 and get \[y-2(y-1)\geq 6\]

OpenStudy (anonymous):

oops \[y-2(y-1)\geq 12\] looks better since \(4\times 3=12\) sorry about that

OpenStudy (anonymous):

you good from there or you need help finishing it?

OpenStudy (anonymous):

(> or equal to) 10

OpenStudy (anonymous):

i doubt it lets check

OpenStudy (anonymous):

remove parentheses and get \[y-2y+2\geq 12\]

OpenStudy (anonymous):

combine like terms \[-y+2\geq 12\]

OpenStudy (anonymous):

subtract \(2\) and get \[-y\geq 10\]

OpenStudy (anonymous):

but you need to solve for \(y\) not \(-y\) change the signs and the inequality to get \[y\leq -10\]